OpenStudy (anonymous):

The drawing bellow shows the dimensions of a tennis court. the width of the singles court is 3/4 of the width of the doubles court. doubles court are 36 feet width and 78 feet length. what is the area of the singles court? a) 2,808 square feet, b) 3,744 square feet, c)1,580 square feet d)2,106 square feet??

OpenStudy (anonymous):

Do I assume that the doubles court and singles court are the same length? I don't play tennis...

OpenStudy (anonymous):

... if so, then 3/4 of 36 = 3/4* 36 = 27 for the width of the singles court, then the length is 78. length times width is the area.

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

so is 2,106

OpenStudy (anonymous):

thank you